The value v of a shopkeeper's inventory software program, in hundred of dollars, is given by v= -3/4t +6,where t is the number of years since the shopowner first bought the program.
Graph the equation and use the graph to estimate what the program is worth 4 years after is was first purchased.

Answer by stanbon(75887)

The value v of a shopkeeper's inventory software program, in hundred of dollars, is given by v= -3/4t +6,where t is the number of years since the shopowner first bought the program.
Graph the equation and use the graph to estimate what the program is worth 4 years after is was first purchased.
--------
Value(4) = (-3/4)4 + 6 = -3+6 = 3
Ans: $300
-------------------
Cheers,
Stan H.
-------------------